# Mixture of Routers

### v2

@@ Abstract
Supervised fine-tuning (SFT) is a milestone in aligning large language models with human instructions and adapting them to downstream tasks. In particular, Low-Rank Adaptation (LoRA) has gained widespread attention due to its parameter efficiency. However, its impact on improving the performance of large models remains limited. Recent studies suggest that combining LoRA with Mixture-of-Experts (MoE) can significantly enhance fine-tuning performance. MoE adapts to the diversity and complexity of datasets by dynamically selecting the most suitable experts, thereby improving task accuracy and efficiency. Despite impressive results, recent studies reveal issues in the MoE routing mechanism, such as incorrect assignments and imbalanced expert allocation. Inspired by the principles of Redundancy and Fault Tolerance Theory. We innovatively integrate the concept of Mixture of Experts into the routing mechanism and propose an efficient fine-tuning method called Mixture of Routers (MoR). It employs multiple sub-routers for joint selection and uses a learnable main router to determine the weights of the sub-routers. The results show that MoR outperforms baseline models on most tasks, achieving an average performance improvement of 1%. MoR can serve as a plug-and-play, parameter-efficient fine-tuning method suitable for a wide range of applications. Our code is available here: https://anonymous.4open.science/r/MoR-DFC6.
Mixture of Routers
Jia-Chen Zhang1, Yu-Jie Xiong1thanks: Corresponding author., Xi-He Qiu1, Chun-Ming Xia1, Fei Dai2, 1School of Electronic and Electrical Engineering, Shanghai University of Engineering Science, 333 Longteng Road, Songjiang District, Shanghai, China 2Fudan University, Key Laboratory of Computational Neuroscience and Brain-Inspired Intelligence, 220 Handan Road, Pudong New District, Shanghai, China Correspondence: xiong@sues.edu.cn
Figure 1: A schematic diagram of the MoR plugin, where the green expert represents the correct expert. MoR effectively corrects misallocations through joint multi-routing assignments.
@@ 1 Introduction
Large Language Models (LLMs) have gradually become the cornerstone of natural language processing (NLP) Devlin et al. (2019); Liu et al. (2021); He et al. (2021); Radford et al. (2019). As model parameters increase, LLMs demonstrate impressive emergent abilities and transfer learning capabilities Wei et al. (2022); Chowdhery et al. (2023); Zhang et al. (2024); Jiang et al. (2024). However, the computational resources required for full fine-tuning are enormous Lv et al. (2024), and more and more research is focusing on parameter-efficient fine-tuning (PEFT) Mangrulkar et al. (2022). The main goal is to significantly reduce the resources required for fine-tuning. For example, P-tuning converts prompts into learnable embedding layers Liu et al. (2022), while LoRA uses a set of low-rank matrices to learn incremental updates Hu et al. (2022). DCFT Zhang et al. (2025) further reduces the number of parameters by deconvolution. Despite the high efficiency of PEFT methods, their fine-tuning performance often falls short of meeting the increasingly complex demands of downstream tasks.
@@
Mixture-of-Experts (MoE) is designed to improve overall model performance by integrating the advantages of multiple expert networks Jiang et al. (2024). The core idea of this approach is that different expert networks can specialize in handling different subsets or features of the data, while a gating mechanism is responsible for determining which expert should process each input Jacobs et al. (1991); Shazeer et al. (2017a). The model typically includes multiple experts and a gating network. The task of the gating network is to evaluate the relevance of each expert based on the input data and dynamically allocate the input to the expert most suited to handle that data. This dynamic allocation mechanism enables MoE models to be more flexible and efficient, especially when dealing with large-scale and diverse datasets Lepikhin et al. (2021); Du et al. (2022); Xue et al. (2021); Zuo et al. (2022). Moreover, the MoE architecture offers higher computational efficiency and scalability. By parallelizing the computation across multiple experts, MoE can optimize resource usage, accelerating both training and inference speed. This is particularly important in scenarios that require processing massive datasets.
@@
Recent research has shown that combining PEFT and MoE allows for the advantages of both methods Zadouri et al. (2024); Liu et al. (2024); Dou et al. (2023). LoRAMoE Dou et al. (2023) is a plugin version of MoE, learns multiple sets of low-rank matrices as experts and uses a softmax layer as a router to compute the input data for each expert. During training, the pre-trained weights are kept frozen, and only the experts and the router are trained. MoLA Gao et al. (2024) further investigates the number of experts at different layers, assigning fewer experts to lower layers and increasing the number of experts as the layer depth increases. While it demonstrates impressive results, recent studies have found issues with the MoE routing mechanism, including incorrect assignments and imbalances in expert allocation Shazeer et al. (2017b); Fedus et al. (2022).
@@
In system design, the principles of Redundancy and Fault Tolerance Theory emphasize the importance of using multiple components to enhance reliability and robustness. By introducing redundancy, systems can mitigate the impact of individual component failures and improve overall performance. Inspired by this theory, we propose a new parameter-efficient MoE method to address the aforementioned issues. Specifically, our approach employs multiple sub-routers for joint decision-making, where each sub-router contributes to the final decision, thereby reducing the risk of errors from any single sub-router. A main router is then used to select the top-r sub-routers based on their scores, ensuring that only the most reliable decisions are prioritized. By adjusting the number of sub-routers, MoR can flexibly adapt to tasks of varying complexity. Ultimately, the weighted cooperation of the sub-routers determines the scores for each expert, and the top-k experts with the highest combined scores are selected for the final inference. We conducted experiments on six benchmarks, including NLP and Commonsense Reasoning (CR) tasks, to demonstrate the effectiveness of MoR. Our main contributions are as follows:
@@
    We conduct numerous experiments on Llama2-7B to validate the effectiveness of MoR. We compare it against benchmarks across six different tasks, and MoR outperforms in most of them.
@@
2 Related Work
When performing SFT tasks, full fine-tuning not only requires substantial computational and storage resources but can also lead to catastrophic forgetting. In contrast, PEFT Mangrulkar et al. (2022) achieves similar results to full fine-tuning by freezing most of the model parameters and training only a small subset of them. Low-Rank Approximation Hu et al. (2022) is a popular and efficient fine-tuning method for LLMs, dubbed as LoRA. It utilizes low-rank approximation theory to effectively adjust the model’s behavior with smaller parameter increments. The forward formula is as follows:
$W=W^{(0)}+\Delta=W^{(0)}+BA,$ (1)
where $\Delta\in\mathbb{R}^{din\times dout}$, $A\in\mathbb{R}^{r\times dout}$, and $B\in\mathbb{R}^{din\times r}$, with $r^{\in(din,dout)}$. The dimensions of $din$ and $dout$ are the same as those of the pre-trained matrix $W$.
Although LoRA significantly reduces the number of parameters, its impact on SFT performance is noticeable. The integration of MoE and LoRA represents a notable trend in recent advancements aimed at improving the performance of LLMs degrees. Zadouri et al. (2024) introduce a novel parameter-efficient MoE framework, Mixture of Vectors (MoV) and Mixture of LoRA (MoLORA), which achieves comparable performance to full fine-tuning with significantly fewer parameters. Huang et al. (2024) propose LoRAHub, which achieves the cross-task generalization capability of LoRA by integrating LoRA modules trained on different tasks through a simple framework. MoLA Gao et al. (2024) experimentally demonstrates that more LoRA experts in higher layers can significantly improve the performance of Transformer-based models.
These methods effectively improved the performance of SFT, but they also led to issues such as errors and imbalanced expert allocation Shazeer et al. (2017b); Fedus et al. (2022). To mitigate the random routing phenomenon observed in MoE, MoELoRA Luo et al. (2024) encourage experts to learn distinct features through contrastive learning, thereby effectively improving model performance. LoRAMoE Dou et al. (2023) integrates multiple LoRA experts through a router. It also mitigates the issue of unbalanced expert utilization via a Localized Balancing Constraint. DeepseekV3 DeepSeek-AI et al. (2024) manages the load through Auxiliary-Loss-Free Load Balancing and Complementary Sequence-Wise Auxiliary Loss. These methods offer valuable insights into errors and imbalanced expert allocation, but our MoR provides a more flexible and plug-and-play strategy for addressing these issues.
@@ 3 Method
In this section, we elaborate on the methodological details of MoR. It is a MoE-style plugin that employs multi-route collaborative expert allocation, as illustrated in Figure 2.
@@ 3.1 Motivation
In this section, we explore the motivation for introducing multiple routers to make joint decisions and explain their necessity and potential advantages based on the theory of redundancy and fault tolerance.
In traditional MoE-style architectures, a single router is responsible for assigning inputs to the most suitable expert network. However, this design has a significant limitation: the single router may make incorrect allocation decisions due to noise, overfitting, or insufficient training, which can lead to a decline in model performance. This vulnerability is particularly pronounced in high-dimensional and complex tasks. The theory of redundancy and fault tolerance provides important inspiration for addressing this issue. The theory suggests that by introducing redundant components and designing appropriate fault-tolerant mechanisms, the reliability and robustness of a system can be significantly improved. Specifically, when one component in the system fails, other components can collaborate to take over its tasks, thereby preventing overall system failure. Inspired by this, we introduce multiple routers in the MoE architecture to make joint decisions. Aggregating multiple independent judgments reduces the bias of a single decision. If one router makes an incorrect judgment due to noise or error, the other routers can correct it through collective decision-making, thereby enhancing the system’s fault tolerance. Moreover, the design of multiple routers introduces the advantage of diversity. Different routers can form complementary decision-making capabilities through varied initializations, training data, or structural designs. This diversity not only strengthens the system’s robustness but also improves the model’s adaptability when facing complex tasks.
3.2 Architecture
@@
where Topk retains the $k$ highest weight distributions P from the Softmax output, and the remaining weights are set to 0. The retained $k$ weights are then re-normalized to ensure the sum of the weights equals 1. In addition, load balancing loss is commonly used in MoE to promote balanced expert selection. To ensure that the original knowledge space of LLMs is not compromised, LoRA is adopted to reduce the occurrence of catastrophic forgetting. By training multiple pairs of low-rank matrices $\{A\}_{i=1}^{N}$ and $\{B\}_{i=1}^{N}$ as experts. The input is then assigned to different experts through a learnable routing module. Our forward equation is shown as follows:
@@
where the main router assigns weights to each of the sub-routers. The final expert routing weight $R^{\prime}$ is a weighted sum of the sub-routers rather than relying on the result of a single router. Specifically, load balancing loss is used in MoR to promote balanced router selection. We summarize the complete training process of MoR in Algorithm 1.
Algorithm 1 MoR Training
1:  Input: Dataset $\mathcal{D}$; total iterations $T$. //Create LoRA experts and routers
2:  Create matrix $N(A,B)$ and main router $W_{R}$ and sub-routers $(W_{r})_{i}$;
3:  for $t=1,\dots,T$ do
4:     Sample a mini-batch from $\mathcal{D}$;//Calculate the score of experts
5:     Compute main router score R as (5);
6:     Compute the expert score $F_{i}(x)^{\prime}$ as (6);
7:     Select the top-k scoring experts;//Iteratively calculate output $y$
8:     for $i=1,\dots,k$ do
9:        $y$ $+=B_{(k)}A_{(k)}(x)$;
10:        return y.
11:     end for
12:  end for
13:  Output: Expert mixture result $y$.
Models SciQA ComQA OpenQA MRPC CoLA RTE Avg.
Full-Fine Tuning 93.12 77.48 80.4 87.13 86.29 87.73 85.36
Prompt Tuning 36.78 37.76 46.2 49.91 59.25 54.17 47.35
LLaMA-Adapter 73.33 73.55 71.8 71.94 47.56 72.93 68.52
LoRA 91.01 75.51 77.0 83.13 86.29 85.92 83.14
LoRAMoE 92.04 78.13 80.0 84.23 86.28 85.20 84.31
MoLA 92.36 78.95 79.6 83.48 86.87 86.28 84.59
LoRAMoE + MoR 92.90 78.54 81.0 84.75 86.39 88.45 85.34
MoLA + MoR 93.08 79.20 82.0 83.94 86.77 88.45 85.57
Table 1: Direct Fine-Tuning Performance comparison of different models on various tasks. We report the accuracy of all the tasks. Higher is better for all metrics. We use the same hyperparameters, which are specified in section C. The best results are denoted in bold.
Method&# Param Metric SciQA ComQA OpenQA MRPC CoLA RTE Avg.
MoLA Acc(%) 92.36 78.95 79.6 83.48 86.87 86.28 84.59
Time(h) 13.20 19.32 9.67 7.52 15.80 5.25 11.79
MoLA + MoR Acc(%) 93.08 79.20 82.0 83.94 86.77 88.45 85.57
$(r=2)$ Time(h) 13.70 19.98 10.03 7.92 17.00 5.55 12.36
MoLA + MoR Acc(%) 93.30 78.05 81.2 84.81 86.10 87.73 85.20
$(r=3)$ Time(h) 13.99 20.33 10.27 7.98 17.15 5.55 12.55
MoLA + MoR Acc(%) 92.99 76.82 80.0 84.00 86.10 88.09 84.67
$(r=4)$ Time(h) 14.15 20.53 10.57 8.02 17.31 5.58 12.69
Table 2: The results of MoR with different router numbers were tested on 6 different tasks, and we report the accuracy and time efficiency. $r$ represents the number of routers. We use the same hyperparameters, with only the router number varying.
@@ 4.1 Experimental Settings
In this section, we design two experimental setups to evaluate the performance of the MoR method, including direct fine-tuning and instruction fine-tuning. Direct fine-tuning refers to fine-tuning the model directly on downstream tasks, while instruction fine-tuning involves first fine-tuning on an instruction-tuning dataset, followed by fine-tuning on the downstream task. LLaMA-2-7B Touvron et al. (2023) is used as the base model. To ensure a fair comparison, LoRAMoE allocates 5 experts per layer, while MoLA adopts a similar progressive expert configuration of 2-4-6-8 as mentioned in its paper. We keep the total number of experts the same across all variants. In both settings, we conduct a grid search on the number of training epochs, considering 10, 15, and 20 epochs for fine-tuning on the downstream task. The highest value from the three experiments is taken as the experimental result. We implement all algorithms using PyTorch Paszke et al. (2019). We use AdamW Loshchilov and Hutter (2017) as the optimizer. We applied LoRA to four weight matrices in the self-attention module $(W_{q},W_{k},W_{v},W_{o})$ and to three weight matrices in the MLP module $(W_{gate},W_{down},W_{up})$. All experiments are conducted on a single NVIDIA A100-80G GPU.
4.1.1 Datasets
For evaluation, we adopt three natural language processing (NLP) tasks and three commonsense reasoning (CR) tasks. For the NLP tasks, we evaluate three popular datasets, including Microsoft’s Research Paraphrase Corpus (MRPC) Dolan and Brockett (2005), Corpus of Linguistic Acceptability (COLA) Wang et al. (2018), and Recognizing Textual Entailment (RTE) Wang et al. (2018). For the CR tasks, we evaluate ScienceQA Lu et al. (2024), CommonsenseQA Talmor et al. (2019), and OpenbookQA Mihaylov et al. (2018).
4.1.2 Baselines
We compare MoR with four popular parameter-efficient tuning methods, including prompt tuning Lester et al. (2021), LLaMAAdapter Zhang et al. (2024), LoRA Hu et al. (2022), LoRAMoE Dou et al. (2023) and MoLA Gao et al. (2024). Additionally, we also evaluate full-parameter fine-tuning.

### v3

@@ Abstract
Supervised fine-tuning (SFT) is a resource-efficient approach for aligning large language models with human instructions. In particular, Low-Rank Adaptation (LoRA) has gained widespread attention due to its parameter efficiency. Recent studies suggest that combining LoRA with Mixture-of-Experts (MoE) significantly enhance fine-tuning performance. MoE enhances task accuracy by dynamically selecting the most suitable experts to adapt to diverse and complex datasets. Despite impressive results, recent studies reveal issues in the MoE routing mechanism, such as incorrect assignments and imbalanced expert allocation. Previous researchers have alleviated this issue by controlling the number of experts or incorporating load balancing losses. However, single-router mechanisms still face challenges such as slow convergence when dealing with a large number of experts. Inspired by redundancy and fault tolerance theory, we are the first to propose the concept of MoE into the routing mechanism and propose an efficient fine-tuning method called Mixture of Routers (MoR). It employs multiple sub-routers for joint selection and uses a learnable main router to determine the weights of the sub-routers. Experimental results show that MoR outperforms baseline models on most tasks, with almost no increase in inference latency. MoR can serve as a plug-and-play, parameter-efficient fine-tuning method for a wide range of applications. Our code is available here: https://github.com/X-Lab-CN/MoR.
keywords:
[…]
@@ 1 Introduction
Large Language Models (LLMs) have gradually become the cornerstone of Natural Language Processing (NLP) Devlin et al. (2019); Radford et al. (2019); Busto-Castiñeira et al. (2025). As model parameters increase, LLMs demonstrate impressive emergent abilities and transfer learning capabilities Wei et al. (2022); Chowdhery et al. (2023); Zhang et al. (2024); Jiang et al. (2024). Therefore, Supervised Fine-Tuning (SFT) gradually became the mainstream solution for addressing various downstream tasks. However, the computational resources required for full fine-tuning were enormous Lv et al. (2024), and more and more research focused on parameter-efficient fine-tuning (PEFT) Mangrulkar et al. (2022). The main goal was to significantly reduce the resources required for fine-tuning. For example, P-tuning converted prompts into learnable embedding layers Liu et al. (2022), while LoRA used a set of low-rank matrices to learn incremental updates Hu et al. (2022). Chain-of-LoRA Qiu et al. (2024) reduced training resource consumption through LoRA networks. Rashid and Amirkhani (2023) achieved unsupervised sentence simplification by fine-tuning BERT. Despite the high efficiency of PEFT methods, their fine-tuning performance often fell short of meeting the increasingly complex demands of downstream tasks.
@@
Mixture-of-Experts (MoE) is designed to improve overall model performance by integrating the advantages of multiple expert networks Jiang et al. (2024). The core idea of this approach is that different expert networks can specialize in handling different subsets or features of the data, while a gating mechanism is responsible for determining which expert should process each input Jacobs et al. (1991); Shazeer et al. (2017b). The model typically includes multiple experts and a gating network. The task of the gating network is to evaluate the relevance of each expert based on the input data and dynamically allocate the input to the expert most suited to handle that data. This dynamic allocation mechanism enables MoE models to be more flexible and efficient, especially when dealing with large-scale and diverse datasets Lepikhin et al. (2021); Du et al. (2022); Xue et al. (2021); Zuo et al. (2022). Moreover, the MoE architecture offers higher computational efficiency and scalability. By parallelizing the computation across multiple experts, MoE can optimize resource usage, accelerating both training and inference speed. This is particularly important in scenarios that require processing massive datasets.
@@
Recent research has shown that combining PEFT and MoE enables the model to leverage the advantages of both approaches Zadouri et al. (2024). LoRAMoE Dou et al. (2023) introduced a plug-in version of MoE that learned multiple sets of low-rank matrices as experts and used a softmax layer as a router to compute each expert’s contribution for a given input. During training, LoRAMoE kept the pre-trained weights frozen and trained only the experts and the router. MoLA Gao et al. (2025) further investigated expert assignment across layers, assigning fewer experts to lower layers and increasing the number of experts with depth. On the other hand, recent studies have identified issues in MoE routing mechanisms, such as incorrect expert assignments and imbalanced expert utilization Shazeer et al. (2017a); Fedus et al. (2022).
@@
In system design, the principles of Redundancy and Fault Tolerance Theory emphasize the importance of using multiple components to enhance reliability and robustness. By introducing redundancy, systems can mitigate the impact of individual component failures and improve overall performance. Inspired by this theory, we propose a new parameter-efficient MoE method to address the aforementioned issues. Our approach employs multiple sub-routers for joint decision-making, where each sub-router contributes to the final decision, thereby reducing the risk of errors from any single sub-router. A main router is used to assign weights to the sub-routers, ensuring that the most reliable decisions are prioritized. Similar to MoE adjusting the number of experts. By adjusting the number of sub-routers, MoR can flexibly adapt to tasks of varying complexity. Ultimately, the weighted cooperation of the sub-routers determines the scores for each expert, and the top-k experts with the highest combined scores are selected for the final inference. Additionally, MoR further balances expert load by introducing load balancing loss. We conducted experiments on six benchmarks, including NLP and Commonsense Reasoning (CR) tasks, to demonstrate the effectiveness of MoR. Our main contributions are as follows:
@@
    The impact of the number of sub-routers on training time and performance is analyzed to facilitate determining the optimal number of sub-routers.
  • $\bullet$
    Numerous experiments are conducted to validate the effectiveness of MoR. We compare it against benchmarks across six different tasks, and MoR outperforms in most of them.
2 Related work
When performing SFT tasks, full fine-tuning not only requires substantial computational and storage resources but can also lead to catastrophic forgetting. In contrast, PEFT Mangrulkar et al. (2022) achieves similar results to full fine-tuning by freezing most of the model parameters and training only a small subset of them. Low-Rank Approximation Hu et al. (2022) is a popular and efficient fine-tuning method for LLMs, dubbed as LoRA. It utilizes low-rank approximation theory to effectively adjust the model’s behavior with smaller parameter increments. Building upon this low-rank adaptation framework, PiSSA Meng et al. (2024) initialized the adapter matrix using the singular values of the original weight matrix and placed the remaining components into a residual matrix. DCFT Zhang et al. (2025a) significantly reduced the number of trainable parameters by completing the incremental update matrix with deconvolution. Bpc-lw Zhang et al. (2025b) fine-tuned the BERT model via LoRA to achieve few-shot HTC tasks. LoRAC Ling et al. (2026) proposed an orthogonal LoRA combination based on QR decomposition, which can preserve previously learned knowledge while integrating new knowledge into continual learning.
Mixture-of-Experts (MoE) architectures enhance model capacity and performance by leveraging multiple specialized expert subnetworks, each trained to handle distinct data patterns or input regions Jiang et al. (2024). A learnable gating network dynamically routes each input to the most relevant expert(s), enabling conditional computation and improved adaptability Shazeer et al. (2017b). This design not only increases model expressiveness but also improves computational efficiency through sparse activation—only a subset of experts is activated per input.
Although LoRA significantly reduces the number of parameters, its impact on SFT performance is noticeable. Zadouri et al. (2024) introduce a novel parameter-efficient MoE framework, MoV and MoLORA, which achieves comparable performance to full fine-tuning with significantly fewer parameters. Huang et al. (2024) propose LoRAHub, which achieves the cross-task generalization capability of LoRA by integrating LoRA modules trained on different tasks through a simple framework. MoLA Gao et al. (2025) experimentally demonstrates that more LoRA experts in higher layers can significantly improve the performance of Transformer-based models.
These methods effectively improved the performance of supervised fine-tuning (SFT), but they also introduced issues such as routing errors and imbalanced expert allocation Shazeer et al. (2017a); Fedus et al. (2022). To mitigate the random routing phenomenon in MoE, MoELoRA Luo et al. (2024) encouraged experts to learn distinct features through contrastive learning, thereby improving model performance. LoRAMoE Dou et al. (2023) integrated multiple LoRA experts via a router and further alleviated unbalanced expert utilization through a Localized Balancing Constraint. DeepseekV3 DeepSeek-AI et al. (2024) addressed load imbalance using Auxiliary-Loss-Free Load Balancing together with Complementary Sequence-Wise Auxiliary Loss. While these approaches provided valuable insights into routing errors and expert imbalance, MoR offers a more flexible and plug-and-play strategy to tackle these challenges.
@@ 3 Method
In this section, we elaborate on the methodological details of MoR. It is a MoE-style plug-in module that employs multi-route collaborative expert allocation, as illustrated in Figure 1.
@@ 3.1 Motivation
In this section, we explore why we introduce multiple routers to make joint decisions and explain their necessity and potential advantages based on the theory of redundancy and fault tolerance.
In traditional MoE-style architectures, a single router is responsible for assigning inputs to the most suitable expert network. However, this design has a significant limitation: the single router may make incorrect allocation decisions due to noise, overfitting, or insufficient training, which can lead to a decline in model performance. This vulnerability is particularly pronounced in high-dimensional and complex tasks. The theory of redundancy and fault tolerance provides important inspiration for addressing this issue. The theory suggests that by introducing redundant components and designing appropriate fault-tolerant mechanisms, the reliability and robustness of a system can be significantly improved. Specifically, when one component in the system fails, other components can collaborate to take over its tasks, thereby preventing overall system failure. Inspired by this, we introduce multiple routers in the MoE architecture to make joint decisions. Different routers can form complementary decision-making capabilities through varied initializations, training data, or structural designs. This diversity not only strengthens the system’s robustness but also improves the model’s adaptability when facing complex tasks.
3.2 Mixture of Routers
@@
where $E_{i}(x)$ denotes the output of the $i$-th expert. Topk retains the $k$ highest weight distributions P from the Softmax output, and the remaining weights are set to 0. The retained $k$ weights are then re-normalized to ensure the sum of the weights equals 1. In addition, load balancing loss is commonly used in MoE to promote balanced expert selection. To ensure that the original knowledge space of LLMs is not compromised, LoRA is adopted to reduce the occurrence of catastrophic forgetting. By training multiple pairs of low-rank matrices $\{A\}_{i=1}^{N}$ and $\{B\}_{i=1}^{N}$ as experts. The input is then assigned to different experts through a learnable routing module. Our forward equation is shown as follows:
@@
where the main router assigns weights to each of the sub-routers. The final expert routing weight $R^{\prime}$ is a weighted sum of the sub-routers rather than relying on the result of a single router. Load balancing loss is employed to promote balanced router selection. It is used to ensure that each expert is utilized reasonably when handling tasks, avoiding situations where some experts are overloaded while others remain idle. By introducing load balancing loss, the model can dynamically adjust the routing mechanism during training, enabling a more even distribution of input samples among different experts. The calculation formula of $L_{\text{balance}}$ is as follows:
$L_{\text{balance}}=\sum_{i=1}^{N_{\text{layers}}}\left(N_{i}\cdot\sum_{j=1}^{N_{i}}t_{j,i}\cdot R_{j,i}\right),$ (6)
where $N_{i}$ is the total number of experts in the $i$-th layer. $t_{j,i}$ is the proportion of tokens assigned to the $j$-th expert in the $i$-th layer. $r_{j,i}$ represents the average routing probability of the $j$-th expert in the $i$-th layer.
@@ 4.1 Experimental settings
Similar to Gao et al. (2025), we design two experimental setups to evaluate the performance of the MoR, including direct fine-tuning and transfer fine-tuning. Direct fine-tuning refers to fine-tuning the model directly on downstream tasks, while transfer fine-tuning involves first performing preliminary fine-tuning on an instruction-tuning dataset (OpenOrca), followed by secondary fine-tuning on the downstream task. The direct fine-tuning experiments use LLaMA-7B-hf Touvron et al. (2023a), LLaMA-2-7B-hf Touvron et al. (2023b), and Mistral-7B-v0.1 Jiang et al. (2023) as the base models. To ensure a fair comparison, LoRAMoE allocates 5 experts per layer, while MoLA adopts a similar progressive expert configuration of 2-4-6-8 as mentioned in its paper. The total number of experts was held constant across all variants. In both settings, we conduct a grid search on the number of training epochs, considering 10, 15, and 20 epochs for fine-tuning on the downstream task. The highest value from the three experiments is taken as the experimental result. We use AdamW Loshchilov and Hutter (2017) as the optimizer. We applied LoRA to four weight matrices in the self-attention module $(W_{q},W_{k},W_{v},W_{o})$ and to three weight matrices in the MLP module $(W_{gate},W_{down},W_{up})$. The MoR results in the main experiments are all based on tests with dual sub-routers. All experiments are conducted on a single NVIDIA A100-80G GPU.
Table 1: Direct Fine-Tuning Performance comparison of different models on various tasks. We report the accuracy of all the tasks. Higher is better for all metrics. The best results are denoted in bold.
Models SciQA ComQA OpenQA MRPC CoLA RTE Avg.
Llama-7B-hf
LoRA 89.12 73.96 75.80 82.43 84.18 83.03 81.42
LoRAMoE 90.28 75.10 76.00 82.43 84.18 83.03 81.84
LoRAMoE + MoR (+0.55) (+0.58) (+0.40) (+2.15) (+1.63) (+0.72) 82.84 (+1.00)
MoLA 90.10 75.42 78.60 83.36 84.64 84.43 82.76
MoLA + MoR (+0.91) (+1.24) (-0.20) (+0.99) (+0.88) (+0.41) 83.46 (+0.70)
Llama-2-7B-hf
LoRA 91.01 75.51 77.00 83.13 86.29 85.92 83.14
LoRAMoE 92.04 78.13 80.00 84.23 86.28 85.20 84.31
LoRAMoE + MoR (+0.86) (+0.41) (+1.00) (+0.62) (+0.11) (+3.25) 85.34 (+1.04)
MoLA 92.36 78.95 79.60 83.48 86.87 86.28 84.59
MoLA + MoR (+0.72) (+0.25) (+2.40) (+0.46) (+0.00) (+2.17) 85.57 (+0.99)
Mistral-7B-v0.1
LoRA 91.97 80.86 87.20 86.83 87.92 89.53 87.39
LoRAMoE 94.46 81.90 88.00 85.73 87.34 88.44 87.65
LoRAMoE + MoR (+0.56) (+1.47) (+0.00) (+1.46) (-0.28) (+1.09) 88.36 (+0.71)
MoLA 94.91 82.96 88.00 86.95 87.24 88.80 88.32
MoLA + MoR (+0.23) (+0.00) (+0.20) (+0.64) (+0.10) (+2.53) 88.76 (+0.44)